# Fair Algorithms with Probing for Multi-Agent Multi-Armed Bandits

### v4

@@ Introduction
Maximizing total expected reward is a common objective in the MA-MAB framework, but it can lead to inequalities in practice (NIPS2016_eb163727). Optimizing for aggregate performance often results in a concentration of profitable arms among a few agents, disadvantaging others (pandp.20181018; pmlr-v32-agarwalb14). This is especially concerning in applications where fair resource assignment is essential, such as in ridesharing platforms, where drivers need equitable access to profitable areas (bubeck2012; Lattimore_Szepesvári_2020), and in content recommendation systems, where creator exposure should not be monopolized (abdollahpouri2020multistakeholder).
@@
Many MA-MAB studies aim to improve fairness, often by altering assignment strategies or adding constraints to naive utilitarian objectives (Meritocratic2018; Patil_Ghalme_Nair_Narahari_2020). A natural baseline is to maximize the sum of agent rewards (utilitarian welfare), but this can produce highly imbalanced assignments: profitable arms tend to be concentrated on a few agents while others receive little or nothing. This effect—where some agents are effectively “starved” of reward despite high aggregate system utility—is well documented in assignment and scheduling literature and is referred to as starvation. Maximizing total reward without regard to its distribution can therefore systematically disadvantage certain agents even when the system as a whole appears efficient. To address this, a line of work replaces or augments the sum objective with fairness-aware criteria such as Nash Social Welfare (NSW), which takes the product (equivalently the geometric mean) of individual utilities, thereby discouraging assignments that leave any agent with very low reward and yielding a balanced equity-efficiency trade-off (NEURIPS2018_be3159ad; Zhang_Conitzer_2021). In particular, jones2023efficient demonstrated that optimizing NSW in MA-MAB settings can prevent persistent exclusion and achieve simultaneous improvements in fairness and overall utility.
@@
However, a common limitation of existing approaches is their dependence on instantaneous reward feedback for updating estimates and steering assignment policies, even though in many real-world scenarios key information is either unobserved or noisy (li2016collaborative). For instance, in ridesharing platforms, uncertainty about passenger demand and fluctuating road conditions can corrupt the platform’s estimates of driver-region rewards. Those corrupted estimates then propagate into the assignment step, potentially producing assignments that are unfair—for example, by repeatedly starving certain drivers who are misestimated as low-reward. To mitigate such risks, probing can be used to actively gather targeted information before making assignments (NEURIPS2019_36d75342). Originating in economics (pandora-Weitzman) as a cost-bearing method to reduce uncertainty in sequential decision-making, probing has been adapted in domains such as database query optimization (deshpande2016approximation; liu2008near), real-time traffic-aware vehicle routing (bhaskara2020adaptive; xu2025online), and wireless network scheduling (abs-2108-03297; xu2023online).
@@
Integrating probing into MA-MAB poses the challenge of balancing exploration and exploitation while maintaining fairness. Related probing problems have been shown to be NP-hard (goel2006asking), and while previous works (Zuo_Zhang_Joe-Wong_2020), have explored probing strategies in MA-MAB, they have simplified assumptions, such as limiting rewards to Bernoulli distributions and ignoring fair assignment. Our framework overcomes these limitations by considering general reward distributions, ensuring fairness, and introducing a probing budget to optimize performance under exploration constraints.
@@
(1) We extend the multi-agent MAB framework with a novel probing mechanism that tests selected arms before assignment. This approach ensures fairness through Nash Social Welfare optimization, departing from previous work that focuses solely on the sum of rewards (Zuo_Zhang_Joe-Wong_2020).
@@
The multi‑armed bandit (MAB) framework has been key for sequential decision‑making under uncertainty (LAI19854; garivier2013; 892067). While early MAB models involve a single decision‑maker choosing one arm per round, many real‑world problems involve multiple agents acting simultaneously (NEURIPS2019_85353d3b), each potentially choosing different arms (NEURIPS2021_c96ebeee). Most existing multi‑agent MAB methods focus on maximizing the total sum of rewards, which can unfairly favor some agents. To address this, researchers have proposed fairness‑aware approaches (NIPS2016_eb163727), with Nash Social Welfare (NSW) (Kaneko1979) proving effective because it maximizes the product of agents’ utilities.
In practice, key aspects of reward distributions are often unknown (8895728; Lattimore_Szepesvári_2020). Existing fair MA‑MAB methods generally rely on passive feedback (Liu_2010; Gai2014DistributedSO). Probing, or “active exploration,” seeks extra information by testing a subset of arms before committing (Chen_Javdani_2015; NIPS2014_66368270), and is valuable when exploring poorly understood arms is risky (golovin_krause_2011; bhaskara2020adaptive).
Recent single‑agent studies formally introduce probing costs: tucker2023costlyobs analyze bandits with costly reward observations, giving matching $\Theta(c^{1/3}T^{2/3})$ bounds; elumar2024costlyprobe allow paying to probe one arm per round and achieve $\tilde{\mathcal{O}}(\sqrt{KT})$ regret; and Observe‑Before‑Play bandits (Zuo_Zhang_Joe-Wong_2020; zhang2020conversational; zuo2022hierarchical) permit a limited number of pre‑observations each round. Offline work has also examined submodular probing for routing problems (bhaskara2020adaptive). However, the bulk of probing research still optimizes aggregate metrics—coverage, latency, or total reward—without incorporating inter‑agent equity. This leaves a gap between fair multi‑agent MAB methods that rely on passive feedback and probing strategies that ignore fairness. We close this gap by coupling cost‑aware, submodular probe selection with NSW‑oriented assignment, so information gained through active exploration translates directly into fair outcomes for all agents.
@@
$\sum_{a}\pi_{j,a,t}=1\quad(\forall\,j),\qquad\sum_{j}\pi_{j,a,t}\leq 1\quad(\forall\,a),$
ensuring, in expectation, one arm per agent and no arm overbooked. As in jones2023efficient, $\pi_{j,a,t}$ may be fractional, reflecting randomized assignment commonly used in online platforms.
@@
We adopt Nash Social Welfare (NSW) as our fairness objective for three key reasons: (i) Pareto efficiency —maximizing the geometric mean never sacrifices total reward when a Pareto‑improvement is possible; (ii) Scale invariance — multiplying all utilities by the same constant leaves the maximizer unchanged, preventing bias due to units of measurement; (iii) Balanced equity‑efficiency trade‑off — NSW penalizes inequality more gently than max–min fairness while still discouraging highly skewed assignments, offering a smooth continuum between purely utilitarian and purely egalitarian objectives (Caragiannis2019; THOMSON2011393).
@@
This formulation accounts for the trade-off between exploration and exploitation. When more arms are probed, the decision-maker obtains more accurate information about reward distributions, leading to better assignment decisions. However, probing incurs costs, such as time delays, energy consumption, or computational overhead, which reduce the net benefit. For instance, in a wireless scheduling scenario, probing more channels provides better estimates of channel conditions but increases latency, reducing the system’s effective throughput (abs-2108-03297). This formulation is appropriate because the probing impact often scales with the system’s overall performance, and similar scaling effects have been modeled in prior work (abs-2108-03297; xu2025online). The function $\alpha(|S_{t}|)$ captures this diminishing return, ensuring that excessive probing is discouraged.
@@
In the offline setting, all reward distributions are known in advance, reducing the problem to a static optimization over the probing set $S$ without the time index $t$. Given any fixed $S$, the optimal assignment policy $\pi$ (cole2015approximating) can be computed, allowing the effective reward to be expressed as a function of $S$.
This optimization is computationally challenging, as similar problems have been shown to be NP-hard (goel2006asking). To address this, we develop a greedy algorithm based on submodular maximization techniques to obtain an approximate solution while accounting for probing costs.
@@
Given a probing set $S$, an assignment policy $\pi=[\pi_{j,a}]$ can be applied to assign arms to agents. For any fixed $S$, one may compute the optimal assignment policy that maximizes the expected Nash Social Welfare (cole2015approximating). Hence, we define the effective objective as
@@
Taking logarithms converts the product $\prod_{j}(\dots)$ into the sum $\log g(S)$. Unfortunately, the resulting set function is still non‑additive—its marginal gain depends on the current value g(S)g(S)—and generally non‑submodular, so classical greedy guarantees no longer apply. Even a much simpler problem is already intractable: selecting at most $k$ items to maximise a strictly concave, increasing function of their total weight is NP‑hard to approximate within any constant factor (AhmedAtamturk2017). Consequently, directly optimising $\log g(S)$ is computationally challenging. To regain tractability, we adopt the classical idea of piecewise‑linear upper envelopes for concave functions(HorstTuyBook; TawarmalaniSahinidisBook). Specifically, we construct a piecewise-linear, concave, and non-decreasing function $\phi:[0,x_{\max}]\to\mathbb{R}$ such that for all $x>0$, we have $\phi(x)\geq\log(x)$. We then define $f_{\mathrm{upper}}(S)=\phi\bigl(g(S)\bigr)$. This construction yields an objective that (i) upper-bounds $\log\bigl(g(S)\bigr)$ and (ii) exhibits diminishing returns, as explained next.
@@
Finally, we can greedily pick arms one by one to maximize the incremental gain in $f_{\text{upper}}(S)$, subject to budget $I$. By standard results on submodular maximization with cardinality constraints (Iyer2013), this yields a (11/e)(1-1/e)-approximation for maximizing $f_{\text{upper}}$. Moreover, we can combine it with the overhead term $(1-\alpha(|S|))$ to trade off between total payoff and probing cost.
@@
We evaluate our framework using controlled simulations and a real-world ridesharing case study. In our simulated environment, we consider a multi-agent multi-armed bandit (MA-MAB) setting with $M$ agents and $A$ arms (e.g., $M=12,A=8$ for small-scale and $M=20,A=10$ for large-scale scenarios). For each agent–arm pair $(j,a)$, rewards are generated as i.i.d. samples from a fixed distribution $D_{j,a}$ with mean $\mu_{j,a}$. In the real-world setting, we apply our framework to the NYYellowTaxi 2016 dataset (shah2020neural), treating vehicles as agents and discretized pickup locations (binned into 0.01° grids) as arms. Rewards are determined by the normalized Manhattan distance between vehicles and pickup points—closer distances yield higher rewards. Vehicle locations are randomly pre-sampled within city bounds and remain fixed, underscoring the practical effectiveness of our approach.
@@
Three algorithms serve as our comparison baselines: Non-Probing, a fair MAB algorithm from jones2023efficient without probing capability, focusing only on optimal assignment with current information; Random Probing with Random assignment (Random P+A), which randomly selects a fixed number of arms for probing and then assigns the arms randomly; and Greedy Probing with Random assignment (Greedy P+A), which uses the same greedy probing strategy as our algorithm but performs random assignment after probing.
@@ Acknowledgments
Fair Algorithms with Probing for Multi-Agent Multi-Armed Bandits
(Supplementary Material)
@@
Bounding the linear part. By (jones2023efficient, Lemma 4.4), there is an event of probability at least $1-\tfrac{\delta}{2}$ on which
@@
To continue the analysis, we need to bound the summation term containing $N_{j,a,t}$. According to a result from previous work, on an event with probability at least $1-\frac{\delta}{2}$, the following inequality holds (jones2023efficient, Lemma 4.4):
@@
Thus, by summing over $t,j,a$ and applying (jones2023efficient, Lemma 4.4) (and Cauchy–Schwarz for the square-root term), one obtains an upper bound of the form
@@
Using the fact that $\hat{\mu}_{j,a,t}-\hat{\mu}_{j,a,t}^{2}\leq 1$ and applying (jones2023efficient, Lemma 4.4) yields

### v5

@@ Introduction
Maximizing total expected reward is a common objective in the MA-MAB framework, but it can lead to inequalities in practice (Joseph et al. 2016). Optimizing for aggregate performance often results in a concentration of profitable arms among a few agents, disadvantaging others (Kleinberg et al. 2018; Agarwal et al. 2014). This is especially concerning in applications where fair resource assignment is essential, such as in ridesharing platforms, where drivers need equitable access to profitable areas (Bubeck et al. 2012; Lattimore_Szepesvári_2020), and in content recommendation systems, where creator exposure should not be monopolized (Abdollahpouri et al. 2020).
@@
Many MA-MAB studies aim to improve fairness, often by altering assignment strategies or adding constraints to naive utilitarian objectives (Joseph et al. 2018; Patil et al. 2021). A natural baseline is to maximize the sum of agent rewards (utilitarian welfare), but this can produce highly imbalanced assignments: profitable arms tend to be concentrated on a few agents while others receive little or nothing. This effect—where some agents are effectively “starved” of reward despite high aggregate system utility—is well documented in assignment and scheduling literature and is referred to as starvation. Maximizing total reward without regard to its distribution can therefore systematically disadvantage certain agents even when the system as a whole appears efficient. To address this, a line of work replaces or augments the sum objective with fairness-aware criteria such as Nash Social Welfare (NSW), which takes the product (equivalently the geometric mean) of individual utilities, thereby discouraging assignments that leave any agent with very low reward and yielding a balanced equity-efficiency trade-off (Heidari et al. 2018; Zhang and Conitzer 2021). In particular, Jones et al. (2023) demonstrated that optimizing NSW in MA-MAB settings can prevent persistent exclusion and achieve simultaneous improvements in fairness and overall utility.
@@
However, a common limitation of existing approaches is their dependence on instantaneous reward feedback for updating estimates and steering assignment policies, even though in many real-world scenarios key information is either unobserved or noisy (Li et al. 2016). For instance, in ridesharing platforms, uncertainty about passenger demand and fluctuating road conditions can corrupt the platform’s estimates of driver-region rewards. Those corrupted estimates then propagate into the assignment step, potentially producing assignments that are unfair—for example, by repeatedly starving certain drivers who are misestimated as low-reward. To mitigate such risks, probing can be used to actively gather targeted information before making assignments (Oh and Iyengar 2019). Originating in economics (Weitzman 1979) as a cost-bearing method to reduce uncertainty in sequential decision-making, probing has been adapted in domains such as database query optimization (Deshpande et al. 2016; Liu et al. 2008), real-time traffic-aware vehicle routing (Bhaskara et al. 2020; Xu et al. 2025), and wireless network scheduling (Xu et al. 2021, 2023).
@@
Integrating probing into MA-MAB poses the challenge of balancing exploration and exploitation while maintaining fairness. Related probing problems have been shown to be NP-hard (Goel et al. 2006), and while previous works (Zuo et al. 2020), have explored probing strategies in MA-MAB, they have simplified assumptions, such as limiting rewards to Bernoulli distributions and ignoring fair assignment. Our framework overcomes these limitations by considering general reward distributions, ensuring fairness, and introducing a probing budget to optimize performance under exploration constraints.
@@
(1) We extend the multi-agent MAB framework with a novel probing mechanism that tests selected arms before assignment. This approach ensures fairness through Nash Social Welfare optimization, departing from previous work that focuses solely on the sum of rewards (Zuo et al. 2020).
@@
The multi‑armed bandit (MAB) framework has been key for sequential decision‑making under uncertainty (Lai and Robbins 1985; Garivier and Cappé 2011; Kumar and Kleinberg 2000). While early MAB models involve a single decision‑maker choosing one arm per round, many real‑world problems involve multiple agents acting simultaneously (Martínez-Rubio et al. 2019), each potentially choosing different arms (Hossain et al. 2021). Most existing multi‑agent MAB methods focus on maximizing the total sum of rewards, which can unfairly favor some agents. To address this, researchers have proposed fairness‑aware approaches (Joseph et al. 2016), with Nash Social Welfare (NSW) (Kaneko and Nakamura 1979) proving effective because it maximizes the product of agents’ utilities.
In practice, key aspects of reward distributions are often unknown (Slivkins 2019; Lattimore_Szepesvári_2020). Existing fair MA‑MAB methods generally rely on passive feedback (Liu and Zhao 2010; Gai and Krishnamachari 2014). Probing, or “active exploration,” seeks extra information by testing a subset of arms before committing (Chen et al. 2015; Amin et al. 2014), and is valuable when exploring poorly understood arms is risky (Golovin and Krause 2011; Bhaskara et al. 2020).
Recent single‑agent studies formally introduce probing costs: Tucker et al. (2023) analyze bandits with costly reward observations, giving matching $\Theta(c^{1/3}T^{2/3})$ bounds; Elumar et al. (2024) allow paying to probe one arm per round and achieve $\tilde{\mathcal{O}}(\sqrt{KT})$ regret; and Observe‑Before‑Play bandits (Zuo et al. 2020; Zhang et al. 2020; Zuo et al. 2022) permit a limited number of pre‑observations each round. Offline work has also examined submodular probing for routing problems (Bhaskara et al. 2020). However, the bulk of probing research still optimizes aggregate metrics—coverage, latency, or total reward—without incorporating inter‑agent equity. This leaves a gap between fair multi‑agent MAB methods that rely on passive feedback and probing strategies that ignore fairness. We close this gap by coupling cost‑aware, submodular probe selection with NSW‑oriented assignment, so information gained through active exploration translates directly into fair outcomes for all agents.
@@
$\sum_{a}\pi_{j,a,t}\leq 1\quad(\forall\,j),\qquad\sum_{j}\pi_{j,a,t}\leq 1\quad(\forall\,a),$
ensuring at most one arm per agent and no arm is overbooked. As in Jones et al. (2023), $\pi_{j,a,t}$ may be fractional, reflecting randomized assignment commonly used in online platforms.
@@
We adopt Nash Social Welfare (NSW) as our fairness objective for three key reasons: (i) Pareto efficiency —maximizing the geometric mean never sacrifices total reward when a Pareto‑improvement is possible; (ii) Scale invariance — multiplying all utilities by the same constant leaves the maximizer unchanged, preventing bias due to units of measurement; (iii) Balanced equity‑efficiency trade‑off — NSW penalizes inequality more gently than max–min fairness while still discouraging highly skewed assignments, offering a smooth continuum between purely utilitarian and purely egalitarian objectives (Caragiannis et al. 2019; Thomson 2011).
@@
This formulation accounts for the trade-off between exploration and exploitation. When more arms are probed, the decision-maker obtains more accurate information about reward distributions, leading to better assignment decisions. However, probing incurs costs, such as time delays, energy consumption, or computational overhead, which reduce the net benefit. For instance, in a wireless scheduling scenario, probing more channels provides better estimates of channel conditions but increases latency, reducing the system’s effective throughput (Xu et al. 2021). This formulation is appropriate because the probing impact often scales with the system’s overall performance, and similar scaling effects have been modeled in prior work (Xu et al. 2021, 2025). The function $\alpha(|S_{t}|)$ captures this diminishing return, ensuring that excessive probing is discouraged.
@@
In the offline setting, all reward distributions are known in advance, reducing the problem to a static optimization over the probing set $S$ without the time index $t$. Given any fixed $S$, the optimal assignment policy $\pi$ (Cole and Gkatzelis 2015) can be computed, allowing the effective reward to be expressed as a function of $S$.
This optimization is computationally challenging, as similar problems have been shown to be NP-hard (Goel et al. 2006). To address this, we develop a greedy algorithm based on submodular maximization techniques to obtain an approximate solution while accounting for probing costs.
@@
Given a probing set $S$, an assignment policy $\pi=[\pi_{j,a}]$ can be applied to assign arms to agents. For any fixed $S$, one may compute the optimal assignment policy that maximizes the expected Nash Social Welfare (Cole and Gkatzelis 2015). Hence, we define the effective objective as
@@
Taking logarithms converts the product $\prod_{j}(\dots)$ into the sum $\log g(S)$. Unfortunately, the resulting set function is still non‑additive—its marginal gain depends on the current value g(S)g(S)—and generally non‑submodular, so classical greedy guarantees no longer apply. Even a much simpler problem is already intractable: selecting at most $k$ items to maximise a strictly concave, increasing function of their total weight is NP‑hard to approximate within any constant factor (Ahmed and Atamtürk 2011). Consequently, directly optimising $\log g(S)$ is computationally challenging. To regain tractability, we adopt the classical idea of piecewise‑linear upper envelopes for concave functions(Salhi 1994; Tawarmalani and Sahinidis 2013). Specifically, we construct a piecewise-linear, concave, and non-decreasing function $\phi:[0,x_{\max}]\to\mathbb{R}$ such that for all $x>0$, we have $\phi(x)\geq\log(x)$. We then define $f_{\mathrm{upper}}(S)=\phi\bigl(g(S)\bigr)$. This construction yields an objective that (i) upper-bounds $\log\bigl(g(S)\bigr)$ and (ii) exhibits diminishing returns, as explained next.
@@
Finally, we can greedily pick arms one by one to maximize the incremental gain in $f_{\text{upper}}(S)$, subject to budget $I$. By standard results on submodular maximization with cardinality constraints (Iyer and Bilmes 2013), this yields a (11/e)(1-1/e)-approximation for maximizing $f_{\text{upper}}$. Moreover, we can combine it with the overhead term $(1-\alpha(|S|))$ to trade off between total payoff and probing cost.
@@
We evaluate our framework using controlled simulations and a real-world ridesharing case study. In our simulated environment, we consider a multi-agent multi-armed bandit (MA-MAB) setting with $M$ agents and $A$ arms (e.g., $M=12,A=8$ for small-scale and $M=20,A=10$ for large-scale scenarios). For each agent–arm pair $(j,a)$, rewards are generated as i.i.d. samples from a fixed distribution $D_{j,a}$ with mean $\mu_{j,a}$. In the real-world setting, we apply our framework to the NYYellowTaxi 2016 dataset (Shah et al. 2020), treating vehicles as agents and discretized pickup locations (binned into 0.01° grids) as arms. Rewards are determined by the normalized Manhattan distance between vehicles and pickup points—closer distances yield higher rewards. Vehicle locations are randomly pre-sampled within city bounds and remain fixed, underscoring the practical effectiveness of our approach.
@@
Three algorithms serve as our comparison baselines: Non-Probing, a fair MAB algorithm from Jones et al. (2023) without probing capability, focusing only on optimal assignment with current information; Random Probing with Random assignment (Random P+A), which randomly selects a fixed number of arms for probing and then assigns the arms randomly; and Greedy Probing with Random assignment (Greedy P+A), which uses the same greedy probing strategy as our algorithm but performs random assignment after probing.
@@ Acknowledgments
Bounding the linear part. By (Jones et al. 2023, Lemma 4.4), there is an event of probability at least $1-\tfrac{\delta}{2}$ on which
@@
To continue the analysis, we need to bound the summation term containing $N_{j,a,t}$. According to a result from previous work, on an event with probability at least $1-\frac{\delta}{2}$, the following inequality holds (Jones et al. 2023, Lemma 4.4):
@@
Thus, by summing over $t,j,a$ and applying (Jones et al. 2023, Lemma 4.4) (and Cauchy–Schwarz for the square-root term), one obtains an upper bound of the form
@@
Using the fact that $\hat{\mu}_{j,a,t}-\hat{\mu}_{j,a,t}^{2}\leq 1$ and applying (Jones et al. 2023, Lemma 4.4) yields